A father is 32 years older than his daughter.In 4 years, the father will be 5 times as old as his daughter. Fid their present ages?
Answer by palanisamy(496)

Let the age of the daughter = x
Then the age of the father = x+32
In 4 years, the father will be 5 times as old as his daughter.
(x+32)+4 = 5(x+4)
x+36=5x+20
36-20=5x-x
16=4x
x=16/4=4
Therefore the age of the daughter = 4
And the age of the father = x+32 = 4+32 = 36